Find 𝑑𝑦

𝑑π‘₯

in each case

(a) 𝑦 = π‘‘π‘Žπ‘›4(3π‘₯)

(3)

(b) π‘₯2𝑦2 + π‘₯𝑠𝑖𝑛𝑦 = 4 use implicit differentiation



1
Expert's answer
2021-10-29T04:54:56-0400

Let us find "\frac{dy}{dx}" in each case

(a) "𝑦 = \tan^4(3𝑥)"

(b) "𝑥^2𝑦^2 + 𝑥\sin𝑦 = 4" use implicit differentiation


(a) "𝑦' = 4\tan^3(3𝑥)(\tan(3𝑥))'= 4\tan^3(3𝑥)\frac{1}{\cos^2 (3x)}(3𝑥)'=
12\frac{\tan^3(3𝑥)}{\cos^2 (3x)}."

(b) Let us use the implicit differentiation:

"2x𝑦^2 + 𝑥^22𝑦y' +\sin𝑦+xy'\cos y = 0."

Therefore, "y'(2𝑥^2y+x\cos y) = -2x𝑦^2 - \sin y," and hence

"y' =- \frac{2x𝑦^2 +\sin y}{2𝑥^2y+x\cos y}."